OpenStudy (hadeel):

∑(n^2) from n= 0 to n =∞ how can i solve it and show that is equal to ∞

OpenStudy (3mar):

Let that helps you.

OpenStudy (hadeel):

actually, i didn't understand

OpenStudy (hadeel):

@3mar

OpenStudy (hadeel):

@mathmale

OpenStudy (sshayer):

\[\sum_{0}^{\infty}n^2=\frac{ n(n+1)(2n+1) }{ 6 }|0 \rightarrow \infty | \rightarrow \infty\] or \[\sum n^2=1^2+2^2+3^2+4^2+...... =1+4+9+16+....\rightarrow \infty~ as~ n \rightarrow \infty\]

OpenStudy (hadeel):

it's very clear now, thanks a lot @3mar @sshayer

OpenStudy (sshayer):

yw

OpenStudy (3mar):

I did not do anything, friend. Thanks a lot.

OpenStudy (3mar):

and thanks for the medal!